OpenStudy (anonymous):

Sandra exclaims that her quadratic with a discriminant of −4 has no real solutions. Sandra then puts down her pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Sandra, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (anonymous):

@Hero

hero (hero):

Hint: \(\sqrt{-1} = i\)

hero (hero):

\(\sqrt{-4} = 4i\)

OpenStudy (anonymous):

is that the formula?

OpenStudy (anonymous):

@Hero

hero (hero):

The quadratic formula is \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) The discriminant is \(b^2 - 4ac\). Notice that the discriminant is under the square root.

hero (hero):

\(i\) represents an imaginary number. You can use \(i\) to represent imaginary solutions whenever you arrive at solutions that are not real.

OpenStudy (anonymous):

but its asking me to create a new question w/ a negative discriminant @Hero

hero (hero):

Yes. I know. I gave you the information you need to help you create a demonstration of how to solve a problem with a negative discriminant. Good luck.

OpenStudy (anonymous):

oh, okay thanks @Hero